# The dark side of the forces: assessing non-conservative force models for atomistic machine learning

### v1

@@ 4.5 Equilibrium properties in the NVT ensemble
The very use of a finite time step in the integration of MD trajectories causes energy fluctuations, and it is not uncommon to use advanced simulation schemes that violate energy conservation (Kühne et al., 2007; Mazzola & Sorella, 2017; Morrone et al., 2011; Laio & Parrinello, 2002), for instance because they use approximations that yield forces contaminated by a stochastic noise. In these cases, using a thermostat can counterbalance the energy error, and obtain stable trajectories that yield configurations and equilibrium average properties close the correct NVT ensemble despite the drift of the conserved quantity that generalizes the total energy for constant-temperature simulations. Judging by the average temperature $\langle T\rangle$ (Table 2), it is relatively easy to control the non-conservative behavior using a white-noise (WN) Langevin thermostat; strong couplings $\tau$ (loosely speaking, the time scale over which the Langevin equation interferes with atomic motion) are needed, as even at 100 fs the average temperature is about 1% off the target value. We discuss how these deviations in the equilibrium temperature affect structural properties of water in Section F.3. The upshot is that for accurate models and strong thermostatting, the effects are small but noticeable. Furthermore, the strong Langevin thermostatting reduces the sampling efficiency, and so even with a respectable 1 ns trajectory, many simple structural averages are not fully converged.
@@ 4.6 Sampling efficiency and time-dependent properties
Aggressive Langevin dynamics is bound to dramatically change time-dependent properties, and in particular to reduce the diffusion coefficient – so that longer trajectories are needed to collect statistically-independent atomic configurations. This slow-down is apparent when looking at the velocity correlation spectra (Figure 3). In the weak coupling regime (WN, $\tau=1000$ fs) there is a (small) increase in diffusion coefficient relative to the reference, because of the unphysically higher temperature, while the high-frequency peaks corresponding to stretching and bending are only weakly perturbed. Stronger couplings alter the dynamics dramatically, and reduce the diffusion coefficient (and hence the efficiency in sampling slow, collective motion) by a factor of about 1.5 ($\tau=100$ fs) and 5 ($\tau=10$ fs), negating the inference speed-up for a non-conservative model – while making it impossible to accurately evaluate any time-dependent property.
@@ 4.7 Geometry optimization
MD performed at very low temperature can be regarded as a form of geometry optimization. Our observations from MD suggest that sufficiently-accurate non-conservative models, combined with good thermostats, should be able to reach reasonable, low-energy structures. We restrict ourselves to optimization algorithms based only on gradients, to avoid the complication of using inconsistent energies, as discussed in Section 3.3, comparing the FIRE (Bitzek et al., 2006) algorithm – that is similar in spirit to zero-temperature MD – and LBFGS (Liu & Nocedal, 1989) – a quasi-Newton algorithm that uses an approximation of the Hessian to accelerate convergence. Comparing different models on the task of optimizing a water snapshot from a MD configuration (Figure 5) shows that inaccurate non-conservative models, such as SOAP-BPNN-NC fail catastrophically at geometry optimization, while more accurate models, such as PET-NC and ORB, can reach a locally-stable configuration, especially using FIRE. We note, however, that non-conservative models are less stable when used with a Hessian-based method, with large fluctuations in the residual force that make it hard to define a stopping criterion. On a practical level, non-conservative forces are bound to make geometry optimization more fragile, and to require careful choice of the minimization algorithm and its convergence parameters, as we observe in Appendix E when comparing different general-purpose models.
@@ 4.8 Non-conservative forces as accelerators
The difficulty of controlling energy conservation when using a direct force model, and the overall instability that we have demonstrated, would indicate that it is better to train conservative models. The idea can, however, be put to good use by combining it with traditional backpropagation-based forces. This is the case of the hybrid PET-M model, which was trained for both types of force evaluation. Such a model can be obtained very efficiently by fine-tuning the energy head of a non-conservative model (see Appendix G).
When using such a “multi-force” model in simulations, one can use the conservative forces for validation, error monitoring and correction, and the direct forces for faster inference. A good example is to use multiple time stepping techniques for molecular dynamics, where the non-conservative forces are used to integrate the equations of motion, and the conservative forces are applied every few steps as a correction. We briefly summarize the logic behind this classic algorithm and demonstrate its application on our MD benchmark in Appendix H. The upshot, as shown in Table 2, Figure 2 and in Section F.3, is that it is possible to obtain results that are effectively indistinguishable from a full conservative evaluation using only 1/8th of the evaluations, recovering almost entirely the speed-up of a non-conservative model, with none of the pathological behaviors. Similar approaches can easily be used to reduce the cost of conservative models in geometry optimization and other applications.
@@ 5 Discussion
Chemical and materials modeling is at the forefront of development in the applications of machine learning to science. The field has long been advocating for the use of physically-informed architectural constraints, but there are indications that its bitter-lesson moment is coming, with the realization that deploying physics-agnostic models at scale provides better outcomes than exploiting physical priors. It appears that this is the case for some of the geometric symmetry constraints (Pozdnyakov & Ceriotti, 2023; Neumann et al., 2024; Qu & Krishnapriyan, 2024), and a growing number of architectures disregard the physical connection between the interatomic potential and the corresponding forces (Gasteiger et al., 2021; Neumann et al., 2024; Liao et al., 2023). With respect to this latter constraint, our study paints a nuanced picture: Atomistic simulations rely on the assumption that forces are the exact derivatives of the potential, and small deviations from this constraints lead to instabilities, even for simple tasks such as the optimization of an atomic structure towards its most stable configuration. Non-conservative behavior also results in molecular dynamics trajectories exhibiting a spontaneous drift away from the desired thermodynamic conditions. Controlling this effect with thermostats requires careful tuning, and disrupts both time-dependent properties and the sampling efficiency of the trajectory, negating the computational advantage of a non-conservative architecture.
@@
Given that the target forces are conservative, accurate models usually exhibit less pronounced non-conservative behavior. As a consequence, one can expect that as the field moves to larger training datasets and more expressive models, some of the pathological effects we observe will become less severe. Our findings, however, suggest that the best way to exploit the speed-up afforded by direct prediction of the forces is not to replace conservative models, but to augment them with a non-conservative head. This can be used to accelerate training by first training a non-conservative model and then fine-tuning its energy head to yield accurate conservative forces through differentiation. The resulting “multi-force” models can also be used to speed up many different types of simulations, by alternating conservative and non-conservative evaluations, avoiding the narrower applicability and inherent instability associated with relying exclusively on direct force predictions.
@@
Code and data required to reproduce the results in this work will be made available upon publication. Molecular dynamics simulations have been performed using the i-PI software (Litman et al., 2024), and geometry optimizations using ASE (Hjorth Larsen et al., 2017). PET models have been trained using the nanoPET implementation from the metatrain library, available at https://github.com/metatensor/metatrain.
@@ Appendix A Forces and potentials as training targets
If we now consider how configurations are sampled at a constant temperature $T$ (running short MD trajecories is a common strategy to generate training sets for atomic-scale models), one sees that each harmonic mode is distributed as $p(q_{k})\propto\exp(-V_{k}(q_{k})/k_{\mathrm{B}}T)=\exp(-m\omega_{k}^{2}q_{k}^{2}/2k_{\mathrm{B}}T)$. As a consequence, one can easily compute the expectation values
@@ Appendix B Theoretical computational cost
However, in the case of backward force evaluation, operations where $x$ is an internal representation and $y$ is a weight can save the $\partial V/\partial y$ calculation. In a simple multi-layer perceptron, where all linear layers correspond to this type of computation, this would yield a backward propagation that is roughly as expensive as the forward pass. This is not the case in transformers, as the attention mechanism involves comparatively expensive operations where both $x$ and $y$ are internal representations, and one can expect the backward propagation of gradients to be somewhere between 1x and 2x the cost of the forward evaluation.
@@ C.1 Timings of general-purpose models
Table 5: Test errors (energies in meV per atom, forces in meV/Å), training and evaluation timings for models trained on the bulk water dataset. Training timings correspond to the timings of a single epoch (in seconds) on 4 H100 GPUs with a total batch size of 64. Evaluation timings correspond to the average timings per atom (in ms) for energy and/or force evaluations for single structures across the test set.
@@ F.3 Structural correlations
Even though the kinetic temperature can be used as a diagnostic of the ensemble error, it does not guarantee that the trajectory yields structural properties consistent with the correct NVT ensemble. We compute different kinds of structural correlation functions: the O-O and H-H pair correlation functions $g_{\mathrm{OO}}(r)$ and $g_{\mathrm{HH}}(r)$ reports on the probability of finding two oxygen or two hydrogen atoms at a given distance $r$, and the orientational correlation function $c_{\mathbf{uu}}(r)$ reports the average scalar product between the orientation vector of two water molecules at a given distance. Results for the PET models (Figure 8) show that structural correlations are consistent for all thermostatting schemes for the conservative runs. The 10 fs white-noise langevin runs exhibit larger error bars, consistent with the poor sampling efficiency. All non-conservative PET-NC runs show statistically-significant discrepancies with respect to the conservative reference. Although it is not possible to rigorously disentangle sampling errors due to non-conservative forces, and the overall difference in the force values between PET and PET-NC, we hypothesize that the differences seen in the strong-Langevin-coupling limit are due to differences in the underlying forcefield, while the differences for the weaker-coupling $\tau=1$ ps and global SVR thermostat runs are associated with the deviation in the sampling temperature.
@@ Appendix G Conservative fine-tuning
Figure 10 shows that a non-conservative model can be trained to its asymptotic accuracy with a 2-4 times smaller investment of computational resources. This is consistent with previous observations (Gasteiger et al., 2021), and one of the main reasons for which direct prediction of the forces has been proposed in previous work. More importantly, the figure also shows a training curve corresponding to a model that has been initially trained with direct $V,\mathbf{f}$ heads (i.e., non-conservatively), and whose $V$ head was then “fine tuned” using energy and back-propagated forces. This “PET-M-FT” model achieves the accuracy of a conservative model trained “from scratch” in about 1/8th of the time, demonstrating that this simple fine-tuning strategy makes it possible to recover most of the speed-up afforded by direct-force training, while having the convenience and accuracy of a conservative model. It might also be possible (although technically more complicated) to alternate gradient steps using only the direct head and steps using the conservative forces to achieve a similar effect.
Figure 10: Training curves (showing validation RMSE force error as a function of the GPU time expenditure on a NVIDIA H100) for a conservative PET-C model, for the non-conservative head of a PET-NC model, and for the fine-tuning of a conservative model initialized from the potential energy head of the PET-NC model.
@@ Appendix H Multiple time stepping
Multiple time stepping (MTS) is based on a Trotter splitting of the Liouville operator (Tuckerman et al., 1992), and can be applied any time one can decompose a potential into a “slowly-varying” and expensive to compute part $V_{\text{slow}}$ and a “quickly-varying”, inexpensive part $V_{\text{fast}}$. The dynamics is propagated using $V_{\text{fast}}$ and a small discrete time step $\Delta t$. Every $M$ steps, one applies a correction based on $V_{\text{slow}}$. In the case of a Velocity-Verlet integrator (Verlet, 1967), a MTS integrator amounts to:

### v2

@@ 4.5 Equilibrium properties in the NVT ensemble
The very use of a finite time step in the integration of MD trajectories causes energy fluctuations, and it is not uncommon to use advanced simulation schemes that violate energy conservation (Kühne et al., 2007; Mazzola & Sorella, 2017; Morrone et al., 2011; Laio & Parrinello, 2002), for instance because they use approximations that yield forces contaminated by a stochastic noise. In these cases, using a thermostat can counterbalance the energy error, and obtain stable trajectories that yield configurations and equilibrium average properties close the correct NVT ensemble despite the drift of the conserved quantity that generalizes the total energy for constant-temperature simulations.
Judging by the average temperature $\langle T\rangle$ (Table 2), it is relatively easy to control the non-conservative behavior using a white-noise (WN) Langevin thermostat. However, strong couplings $\tau$ (the time scale over which the thermostat interferes with atomic motion) are needed, as even at 100 fs the average temperature is significantly off the target value. We discuss how these deviations in the equilibrium temperature affect structural properties of water in Section F.3. The upshot is that for accurate models and strong thermostatting, the effects are small but noticeable. Furthermore, the strong Langevin thermostatting reduces the sampling efficiency, and so even with a respectable 1 ns trajectory, many simple structural averages are not fully converged.
@@ 4.6 Sampling efficiency and time-dependent properties
Aggressive Langevin dynamics is bound to dramatically change time-dependent properties, and in particular to reduce the diffusion coefficient – so that longer trajectories are needed to collect statistically independent atomic configurations. This slow-down is apparent when looking at the velocity correlation spectra (Figure 3). In the weak coupling regime (WN, $\tau=1000$ fs) there is a (small) increase in diffusion coefficient relative to the reference, because of the unphysically higher temperature, while the high-frequency peaks corresponding to stretching and bending are only weakly perturbed. Stronger couplings alter the dynamics dramatically, and reduce the diffusion coefficient (and hence the efficiency in sampling slow, collective motion) by a factor of about 1.5 ($\tau=100$ fs) and 5 ($\tau=10$ fs), negating the inference speed-up for a non-conservative model – while making it impossible to accurately evaluate any time-dependent property.
@@ 4.7 Geometry optimization
MD performed at very low temperature can be regarded as a form of geometry optimization. Our observations from MD suggest that sufficiently accurate non-conservative models, should be able to reach reasonable, low-energy structures. We restrict ourselves to optimization algorithms based only on gradients, to avoid the complication of using inconsistent energies, as discussed in Section 3.3, comparing the FIRE (Bitzek et al., 2006) algorithm – that is similar in spirit to zero-temperature MD – and LBFGS (Liu & Nocedal, 1989) – a quasi-Newton algorithm that uses an approximation of the Hessian to accelerate convergence. Comparing different models on the task of optimizing a water snapshot from a MD configuration (Figure 6) shows that inaccurate non-conservative models, such as SOAP-BPNN-NC, fail catastrophically at geometry optimization, while more accurate models, such as PET-NC and ORB, can reach a locally stable configuration, especially using FIRE. We note, however, that non-conservative models are less stable when used with a Hessian-based method, with large fluctuations in the residual force that make it hard to define a stopping criterion. On a practical level, non-conservative forces are bound to make geometry optimization more fragile, and to require careful choice of the minimization algorithm and its convergence parameters, as we observe in Appendix E when comparing different general-purpose models.
@@ 4.8 Non-conservative forces as accelerators
While we have observed that conservative MLIPs are better suited for practical simulations, we suggest that hybrid models, which additionally support direct, non-conservative, force predictions, can be used for faster inference and training. Such models can be obtained by training both force heads jointly, as demonstrated in the PET-M model, or, more efficiently, by first training a non-conservative model and then fine-tuning its energy head to yield conservative forces. As shown in Figure 4 and further discussed in Appendix G, this leads to the accuracy and physical correctness of conservative models at highly reduced training time.
Figure 4: Training curves for PET-C and PET-NC models, and the fine-tuning of a hybrid PET-M model initialized (epoch marked with arrow) from the potential energy head of the PET-NC model.
In simulations, one can then use the conservative forces of such a hybrid model for validation, error monitoring and correction, and the direct forces for faster inference. A good example is to use multiple time stepping techniques for molecular dynamics, where the non-conservative forces are used to integrate the equations of motion, and the conservative forces are applied every few steps as a correction. This is further explained in Appendix H; results using this technique are shown in Table 2, Figure 2 and in Appendix H and demonstrate that 1/8th of conservative force evaluations are sufficient to obtain results that are essentially indistinguishable from fully conservative ones.
@@ 5 Discussion
Chemical and materials modeling is at the forefront of development in the applications of machine learning to science. The field has long been advocating for the use of physically informed architectural constraints, but there are indications that its bitter-lesson moment is coming, with the realization that deploying physics-agnostic models at scale provides better outcomes than exploiting physical priors. It appears that this is the case for some of the geometric symmetry constraints (Pozdnyakov & Ceriotti, 2023; Neumann et al., 2024; Qu & Krishnapriyan, 2024), and a growing number of architectures disregard the physical connection between the interatomic potential and the corresponding forces (Gasteiger et al., 2021; Neumann et al., 2024; Liao et al., 2023). With respect to this latter constraint, our study paints a nuanced picture. Atomistic simulations rely on the assumption that forces are the exact derivatives of the potential, and small deviations from this constraints lead to instabilities. Non-conservative behavior also results in molecular dynamics trajectories exhibiting a spontaneous drift away from the desired thermodynamic conditions. Controlling this effect with thermostats requires careful tuning, and disrupts both time-dependent properties and the sampling efficiency of the trajectory, negating the computational advantage of a direct-force architecture.
@@
Given that the target forces are conservative, accurate models usually exhibit less pronounced non-conservative behavior. As a consequence, one can expect that, as the field moves to larger training datasets and more expressive models, some of the pathological effects we observe will become less severe. Our findings, however, suggest that the best way to exploit the speed-up afforded by direct prediction of the forces is not to replace conservative models, but to augment them with a non-conservative head. This can be used to accelerate training by first training a non-conservative model and then fine-tuning its energy head to yield accurate conservative forces through differentiation. The resulting “multi-force” models can also be used to speed up many different types of simulations, by alternating conservative and non-conservative evaluations, avoiding the narrower applicability and inherent instability associated with relying exclusively on direct force predictions. This insight enables the efficient training of the next generation of universal machine-learning interatomic potentials while retaining the physical correctness required for practical simulations.
@@
Code and data required to reproduce the results in this work are available on Zenodo at https://zenodo.org/records/14778891. More details on the software used are available in Appendix I.
Impact Statement
This paper presents work whose goal is to advance the application of machine learning to simulations of molecules and materials, which can be used for many purposes. As potential impacts are hard to predict and wide-ranging, we do not believe it is necessary to discuss them here in detail.
@@ Appendix A Forces and potentials as training targets
If we now consider how configurations are sampled at a constant temperature $T$ (running short MD trajectories is a common strategy to generate training sets for MLIPs), one sees that each harmonic mode is distributed as $p(q_{k})\propto\exp(-V_{k}(q_{k})/k_{\mathrm{B}}T)=\exp(-m\omega_{k}^{2}q_{k}^{2}/2k_{\mathrm{B}}T)$. As a consequence, one can easily compute the expectation values
@@ Appendix B Theoretical computational cost
However, in the case of backward force evaluation, operations where $x$ is an internal representation and $y$ is a weight can save the $\partial V/\partial y$ calculation. In a simple multi-layer perceptron, where all linear layers correspond to this type of computation, this would yield a backward propagation that is roughly as expensive as the forward pass. This is not the case in transformers, as the attention mechanism involves comparatively expensive operations where both $x$ and $y$ are internal representations, and one can expect the backward propagation of gradients to be somewhere between 1×1$\times$1 × and 2×2$\times$2 × the cost of the forward evaluation.
@@ C.1 Timings of general-purpose models
Table 5: Test errors (energies in meV per atom, forces in meV/Å), training and evaluation timings for models trained on the bulk water dataset. Training timings correspond to the time to compute a single epoch (in seconds) on 4 H100 GPUs with a total batch size of 64. Evaluation timings correspond to the average time per atom (in ms) for energy and/or force evaluations for single structures across the test set.
@@ F.3 Structural correlations
Even though the kinetic temperature can be used as a diagnostic of the ensemble error, it does not guarantee that the trajectory yields structural properties consistent with the correct NVT ensemble. We compute different kinds of structural correlation functions: the O-O and H-H pair correlation functions $g_{\mathrm{OO}}(r)$ and $g_{\mathrm{HH}}(r)$ reports on the probability of finding two oxygen or two hydrogen atoms at a given distance $r$, and the orientational correlation function $c_{\mathbf{uu}}(r)$ reports the average scalar product between the orientation vector of two water molecules at a given distance. Results for the PET models (Figure 9) show that structural correlations are consistent for all thermostatting schemes for the conservative runs. The 10 fs white-noise langevin runs exhibit larger error bars, consistent with the poor sampling efficiency. All non-conservative PET-NC runs show statistically significant discrepancies with respect to the conservative reference. Although it is not possible to rigorously disentangle sampling errors due to non-conservative forces, and the overall difference in the force values between PET and PET-NC, we hypothesize that the differences seen in the strong-Langevin-coupling limit are due to differences in the underlying forcefield, while the differences for the weaker-coupling $\tau=1$ ps and global SVR thermostat runs are associated with the deviation in the sampling temperature.
@@ Appendix G Conservative fine-tuning
Figure 11 and Figure 12 show that a non-conservative model can be trained to its asymptotic accuracy with a 2-4 times smaller investment of computational resources. This is consistent with previous observations (Gasteiger et al., 2021), and one of the main reasons for which direct prediction of the forces has been proposed in previous work. More importantly, the figures also show training curves corresponding to a model that has been initially trained with direct $V,\mathbf{f}$ heads (i.e., non-conservatively), and whose $V$ head was then “fine tuned” using energy and back-propagated forces (while the non-conservative force head was also trained to avoid its degradation). This “PET-M-FT” model achieves the accuracy of a conservative model trained “from scratch” in about 1/3rd of the total training time, demonstrating that this simple fine-tuning strategy makes it possible to recover most of the speed-up afforded by direct-force training, while having the convenience and accuracy of a conservative model. It might also be possible (although technically more complicated) to alternate gradient steps using only the direct head and steps using the conservative forces to achieve a similar effect.
Figure 11: Training curves (showing validation MAE force error as a function of the GPU time expenditure on a NVIDIA H100) for a conservative PET-C model, for the non-conservative head of a PET-NC model, and for the fine-tuning of a conservative model initialized from the potential energy head of the PET-NC model. Training was conducted on the liquid water dataset.
Figure 12: Training curves (showing validation MAE force error as a function of the GPU time expenditure on a NVIDIA H100) for a conservative PET-C model, for the non-conservative head of a PET-NC model, and for the fine-tuning of a conservative model initialized from the potential energy head of the PET-NC model. Training was conducted on a small subset of OC20 (Chanussot et al., 2021) (training on the first 15000 samples of the 200k S2EF training set, and using the next 5000 for validation).
@@ Appendix H Multiple time stepping
Multiple time stepping (MTS) is based on a Trotter splitting of the Liouville operator (Tuckerman et al., 1992), and can be applied any time one can decompose a potential into a “slowly varying” and expensive to compute part $V_{\text{slow}}$ and a “quickly varying”, inexpensive part $V_{\text{fast}}$. The dynamics is propagated using $V_{\text{fast}}$ and a small discrete time step $\Delta t$. Every $M$ steps, one applies a correction based on $V_{\text{slow}}$. In the case of a Velocity-Verlet integrator (Verlet, 1967), a MTS integrator amounts to:
@@
Appendix I Software
Molecular dynamics simulations have been performed using the i-PI software (Litman et al., 2024), and geometry optimizations using ASE (Hjorth Larsen et al., 2017).